Given y equals ​f(u) and u equals ​g(x), find StartFraction dy Over dx EndFraction equals f prime left parenthesis g left parenthesis x right parenthesis for the following functions y equals sin, Equals 7x plus 4
Answer #1


Given RB (0-7) 3= f(u) U-96) f$95x3) dx y = Sincax) + 4 COSE da dy - coscox) da 902)=22 9023 I foul-sino) Desing flo): cosco)

Please do like